OpenStudy (anonymous):

HElp !! Find the square roots of -1+i

OpenStudy (cggurumanjunath):

IS IT  ?

OpenStudy (anonymous):

thats it

OpenStudy (cggurumanjunath):

THIS SHOULD HELP !

OpenStudy (loser66):

I think, |x + yi| = \(\sqrt{x^2 + y^2}\) yours is |-1 +i | =\(\sqrt{(-1)^2 + 1^2}= \sqrt{2}\) and then \(\sqrt{-1+i}= \sqrt{\sqrt{2}}= 2^{\dfrac{1}{4}}\)\)